Consider two lines L1 and L2 whose direction cosines l1,m1,n1 and l2,m2,n2 are
given by the equations for l,m,n :
al +bm+cn = 0, f mn+gnl +hlm = 0,
where abc 6= 0. Show that if L1 ⊥ L2, then f/a + g/b +h/c = 0.
1
Expert's answer
2020-03-10T10:43:53-0400


As per the given question,


The equation of the lines are

"al+bm+cn=0\u2212\u2212\u2212\u2212\u2212\u2212\u2212(i)"

And

"fmn+gnl+hlm=0\u2212\u2212\u2212\u2212\u2212\u2212\u2212(ii)"

from the equation (i)

"n=-(\dfrac{al+bm}{c})" and "n=-(\dfrac{al+cn}{b})"

Now, substituting the value of n in the equation (ii)


"fm(\dfrac{−(al+bm)}{c})+gl(\dfrac{−(al+bm)}{c})+hlm=0"

"fm(-\dfrac{al+bm}{c})+gl(-\dfrac{al+bm}{c})+hlm=0"

multiplying both side by c

"-fm(al+bm\u200b)-gl(al+bm\u200b)+chlm=0"

now, multiplying both side by -1

"fm(al+bm)+gl(al+bm)\u2212hlmc=0"

"fmal+fbm^2+agl^2+bglm\u2212hlmc=0"


"agl^2+afml+bfm^2+bmgl\u2212hlmc=0"

Now, both side divided by "m^2"

"ag(\dfrac{l}{m})2+(af+bg−ch)(\dfrac{l}{m})+bf=0−−−−−−−−(iii)"

Similarly, after substituting the value of m in the question (ii) and we will rearrange it accordingly,

"ah(\dfrac{l}{n})^2+(af+gb−hc)(\dfrac{l}{n})+fc=0−−−−−−−−(iv)"

"l_1\u200b,m_1\u200b,n_1\u200b" and "l_2,m_2,n_2" are the direction cosine of straight line

roots of the equation (iii)

so, from the equation (iii)

"\dfrac{l_1 l_2}{m_1 m_2}=\dfrac{\dfrac{f}{a}}{\dfrac{g}{b}}=\dfrac{fb}{ag}" similarly, from equation (iv) "\dfrac{l_1 l_2}{n_1 n_2}=\dfrac{fc}{ah}"

So, we can write it as

"\dfrac{l_1 l_2}{\dfrac{f}{a}}=\dfrac{m_1 m_2}{\dfrac{g}{b}}=\dfrac{n_1 n_2}{\dfrac{h}{c}}=k," let it is equal to a constant K

So,

"l_1 l_2+m_1m_2+n_1n_2=0"

"K(\dfrac{f}{a}+\dfrac{g}{b}+\dfrac{h}{c})=0"

Hence, dividing both side by K

"(\dfrac{f}{a}+\dfrac{g}{b}+\dfrac{h}{c})=0"

Hence L1 and L2 are perpendicular.